OBSERVATION OF NEW INFRARED TRANSITIONS IN SOLID PARAHYDROGEN

M. MENGEL, \underline{B. P. WINNEWISSER} AND M. WINNEWISSER, Physikalisch-Chemisches Institut, Justus-Liebig-Universität, Heinrich-Buff-Ring 58, D-35392 Giessen, Germany.

In recent measurements of infrared transitions in solid parahydrogen with low ortho impurities, using an internal multireflection system^{a} for a sample path length of 14 cm, we have detected new weak features in the spectral region 10 000–13 000 cm^{−1}. Of particular interest here is the second overtone band of solid hydrogen and double transitions of the type Q_{2}(J′) + Q_{1}(J) (J, J′ = 0, 1) which will be presented. At 10 241.07 cm^{−1} the new single transition W_{2}(0) could be observed. We will also present further results of the analysis of our measurements in the first overtone region, especially a theoretical explanation of the intensities in the fine structure of the Q_{1}(1) + Q_{1}(1) double transition. The absorption spectrum in the mid-infrared region (700–2 000 cm^{−1}) is currently under investigation, and new information in this region will also be presented.

aR. A. Steinhoff, B. P. Winnewisser and M. Winnewisser, Phys. Rev. Lett. 73, 2833-2836 (1994).